### v1

# Entanglement cones and horizons in analogue cosmological production of Dirac fermions

Abstract
Although gravitational particle production in curved quantum field theories (cQFTs) is key to our understanding of the early universe and black hole physics, its direct observation requires extreme conditions or unrealistic sensitivities. Recent progress in quantum simulators indicates that analogues of gravitational particle production can be observed in table-top experiments of cold atomic gases described by effective cQFTs. This promises a high degree of tunability in the synthesised curved spacetimes and, moreover, sets a clear roadmap to explore the interplay of gravitational particle production with other non-perturbative effects genuine to interacting QFTs. We hereby focus on the appearance of fermion condensates for self-interacting Dirac fermions, and study how dynamical mass generation and spontaneous symmetry breaking affects real-time dynamics. We use the entanglement contour as a tool to analyze the spatio-temporal causal structure of particle production, showing how it can account for the cosmological horizon in accelerating spacetimes, while also being sensitive to the effect of different symmetry-breaking processes. In particular, we show that the combined breakdown of time-reversal symmetry due to the expanding spacetime, and parity due to a pseudo-scalar condensate, manifest through the structure of the light-cone-like propagation of entanglement in this analogue cQFT.
preprint: APS/123-QED
Introduction.– Quantum simulators (QSs) promise to become a transformative tool to explore the complex quantum many-body frontier common to various disciplines of contemporary science  [1, 2, 3]. In theoretical physics, QSs can sometimes realize analogues of phenomena that would otherwise be difficult to observe, if not altogether impossible, in their original realm. These two aspects come together in the study of quantum field theories (QFTs) in curved spacetimes [4, 5, 6, 7, 8, 9, 10] by means of experimental quantum devices [11, 12, 13, 14, 15, 16] that can be used to explore foundational questions about gravity, or test our current cosmological models of the universe. These QSs foster a genuinely multidisciplinary approach that integrates concepts from atomic physics, quantum optics and quantum information theory, with those of high-energy physics and general relativity. A particularly-relevant example is that of quantum entanglement [17, 18, 19], which provides insights into the role of quantum correlations in the complexity frontier of many-body systems, and plays an important role in questions as diverse as the nature of the QFT vacuum or the evaporation of black holes, and also possibly in the information paradox for the structure of spacetime at the quantum level.
In this Letter, we focus on QSs of ultra-cold atoms in Raman optical lattices [20, 21, 22, 23, 24, 25, 26, 27] as analogues of particle production of self-interacting Dirac fermions in an effective spacetime with a tunable accelerated expansion. We demonstrate how the entanglement contour [28], when combined with a non-perturbative variational ansatz based on fermionic Gaussian states [29, 30], can provide a clean spatio-temporal resolution for the real-time dynamics of these Dirac fields. Looking at particle production through this entanglement-contour lens, we unveil interesting phenomena like a cosmological particle horizon in the entanglement spreading, and the impact on the entanglement-cone structure of symmetry-breaking fermion condensates during particle production, pointing at future research opportunities in analogue gravity [31].
The model. We consider a Friedmann-Robertson-Walker spacetime $\mathord{\rm d}s^{2}=\d{d}t^{2}-\mathsf{a}^{2}(t)\mathord{\rm d}\Sigma^{2}$ with a scale factor $\mathsf{a}(t)$. In a cosmological scenario [32], the scale factor connects an inflationary epoch to our current dark-energy and dark-matter dominated universe in $D=3+1$ dimensions. In contrast, in an analogue setup, we are free from these constraints, and can: (i) connect two asymptotic Minkowski regions $\mathsf{a}_{0}\to\sf{a}_{f}$ by an exponential scale factor controlled by the Hubble rate $\mathsf{H}$, such that the duration is $\Delta t=\frac{1}{\mathsf{H}}\log\frac{\mathsf{a}_{\rm f}}{\mathsf{a}_{0}}$; (ii) vary the Hubble rate at will, as it is not fixed by cosmological observations, interpolating between adiabatic $\sf{H}\rightarrow 0$ and sudden $\sf{H}\rightarrow\infty$ expansions; and (iii) explore reduced dimensionalities that may lead to simpler analogue experiments. This flexibility extends to the (iv) dynamical quantum fields that fill in this expanding spacetime which, in the standard cosmological model of inflation, correspond to a scalar field under certain self-interaction potentials. For the reasons exposed below, we shall be interested instead in self-interacting Dirac spinors $\hat{\psi}(x)\to\hat{\psi}_{i}(t)$ discretized on a co-moving chain ${x}_{i}=ia$, where $a$ is the lattice spacing and $i\in\mathbb{Z}_{N_{\rm s}}$. By using conformal time via $\eta=\int{\rm d}t/\mathsf{a}(t)$, and rescaling the fields as $\hat{\psi}_{i}\rightarrow\sqrt{\sf{a}(\eta)}\hat{\psi}_{i}$, there is not need to simulate an inflating lattice with increasing proper distance $d_{ij}(t)=\mathsf{a}(t)|x_{i}-x_{j}|$, and the regularised action reduces to a fermionic QFT with explicit dependence on conformal time. Finally, we can also (iv) explore the subsequent evolution of the produced particles within a flat spacetime $\mathsf{a}_{\rm f}$, which will allow us to derive a neat picture of thermalization as these particles fill in the universe.
We work with a Hamiltonian lattice field theory with rescaled spinors and conformal time associated to the original cQFT action that includes vierbein fields and the spin connection [33]. In particular, we employ a Wilson-type regularization to deal with fermion doubling [34] which, in momentum space, yields a free part $\hat{H}=\sum_{\textrm{k}\in\rm BZ}\hat{\psi}_{\textrm{k}}^{\dagger}h_{\textrm{k}}\big{(}m\mathsf{a}(\eta)\big{)}\hat{\psi}_{\textrm{k}}$ with
$h_{\textrm{k}}\big{(}m\mathsf{a}(\eta)\big{)}=-\frac{\sin\textrm{k}a}{a}\gamma^{0}\gamma^{1}+\left(m\mathsf{a}(\eta)+\frac{1-\cos\textrm{k}a}{a}\right)\gamma^{0},$ (1)
where the matrices are $\gamma^{0}=\sigma^{z}$, $\gamma^{1}={\rm i}\sigma^{y}$ and $\gamma^{5}=\gamma^{0}\gamma^{1}$. This free field theory has a simple interpretation as a tight-binding model on a synthetic ladder, and can be realised with minor modifications of recent experiments with alkaline atoms in optical Raman lattices [27]. Here, the ladder legs represent the spinor components, encoded in two hyperfine levels of ${}^{87}{\rm Sr}$, and an energy imbalance that accounts for the time-dependent mass $m\mathsf{a}(\eta)$ can be controlled by the Raman lasers detuning.
Being quadratic, Wick’s theorem holds for this part, and the real-time dynamics can be fully characterised through the correlation matrix $\Gamma_{ij}(\eta)=\bra{\psi(\eta)}\hat{\psi}^{\phantom{\dagger}}_{i}\hat{\psi}_{j}^{\dagger}\ket{\psi(\eta)}$. Regarding the full many-body Hilbert space, the dynamics is constrained to a reduced manifold of so-called fermionic Gaussian states (fGS) {|$\{\ket{\psi(\Gamma)}:\,\Gamma\in\textrm{GL}_{4N_{\rm s}\!}(\mathbb{C})\}$ during the expansion. This can then be used as a variational family to approximate genuinely-interacting QFTs. In this limit, particle production [4, 5] can be formalised by a Bogoliubov transformation [35, 36] that acts by mapping |$\ket{\psi(\Gamma_{0})}\mapsto\ket{\psi(\Gamma_{\rm f})}$ producing particles. Taking advantage of the above points (i)-(iv), we can actually go beyond this simple limit and explore non-perturbative effects by including self-interactions according to the Gross-Neveu model [37], namely
$\hat{{V}}_{\rm int}=\frac{g_{0}^{2}}{2}\sum_{i}\left(\hat{\psi}_{i}^{\dagger}\gamma^{0}\hat{\psi}^{\phantom{\dagger}}_{i}\right)^{\!\!2}\!\!,$ (2)
where $g_{0}^{2}$ is the dimensionless coupling strength which, in contrast to to the bare mass, does not get renormalised by the scale factor. In a previous work [38], we have indeed shown that the fGS ansatz can be used to approximate the vacuum |$\ket{\psi(\Gamma_{0})}$, and that it coincides exactly with large-$N$ predictions [39]. The fGS predictions account for the onset of scalar and pseudo-scalar fermion condensates, which are specific elements of the correlation matrix of fermion bilinears $\Sigma={g_{0}^{2}}\sum_{\textrm{k}}\langle\hat{\psi}_{\textrm{k}}^{\dagger}\gamma^{0}\hat{\psi}_{\textrm{k}}\rangle/{2aN_{\rm s}}$, $\Pi={{\rm i}g_{0}^{2}}\sum_{\textrm{k}}\langle\hat{\psi}_{\textrm{k}}^{\dagger}\gamma^{1}\hat{\psi}_{\textrm{k}}\rangle/{2aN_{\rm s}}$. During the expansion, these condensates will evolve non-perturbatively affecting the particle production and vice versa, which we can approximate variationally via our fGS, leading to a set of self-consistent non-linear differential equations
${\rm i}\frac{\mathord{\rm d}}{\mathord{\rm d}\eta}\Gamma(\eta)=\left[\tilde{h}_{\textrm{k}}\big{(}m\mathsf{a}(\eta),\Gamma(\eta)\big{)},\;\Gamma(\eta)\right],$ (3)
where $h_{\textrm{k}}\to\tilde{h}_{\textrm{k}}\big{(}m\mathsf{a}(\eta),\Gamma\big{)}={h}_{\textrm{k}}\big{(}m\mathsf{a}(\eta)+\Sigma(\Gamma)\big{)}-\mathord{\rm i}\Pi(\Gamma)\gamma^{1}.$ We see how the non-linearities arise through $\tilde{h}_{\rm k}$ via the self-energy expressed in terms of the condensates, which follows from Wick’s theorem on the fGS manifold $\bra{\psi(\Gamma)}\hat{{V}}_{\rm int}\ket{\psi(\Gamma)}$.
Figure 1: (a) Spatio-temporal propagation of EC in the asymptotic flat region $\mathsf{a}_{\rm f}$, both in the free $g_{0}^{2}=0$ and in the interacting $g_{0}^{2}=1$ regimes, and setting $ma=-1$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, ${\sf H}a=100$ (sudden expansion). The dashed white lines indicate the light-cone-like predictions based on the group velocity after the sudden expansion. (b,c) Spatio-temporal propagation of EC during a de Sitter expansion, using conformal and cosmological time respectively, and setting $ma=1$, ${\sf H}a=0.1$, $g_{0}^{2}=2$, $\mathsf{a}_{0}=1/3$, with $\eta$ and $t$ expressed in lattice units. The expansion in (b) starts at $\eta_{0}a=-30$ and extends to the distant future $\eta\to 0^{-}$, approximated by $\eta a=-0.001362$. The shaded regions denote the asymptotic spatial boundary that particles cannot cross. In (c) the corresponding expansion covers $t_{0}=0$ to $ta\to\infty$, here approximated by $t=100/a$.
Solving these equations with the initial condition |$\ket{\psi(\Gamma_{0})}$ one obtains a time-evolved fGS |$\ket{\psi(\Gamma(\eta_{\rm f}))}$ with a non-vanishing density of particles with respect to the instantaneous groundstate |$\ket{\psi(\Gamma_{\rm f})}$ at $(m\mathsf{a}_{\rm f},g_{0}^{2})$. Within our variational approximation, the interpretation in terms of a Bogoliubov transformation is still useful, but it now depends on the full non-perturbative time history $\{\Gamma\}_{\eta}=\{\Gamma(\eta):\,\eta\in[\eta_{0},\eta_{\rm f}]\}$, allowing to express the evolved state as
$|\psi(\Gamma(\eta_{\rm f}))\rangle=\bigotimes_{\textrm{k}\in\text{BZ}}\left(\alpha_{\textrm{k}}(\{\Gamma\}_{\eta})\ -\beta_{\textrm{k}}(\{\Gamma\}_{\eta})\ a_{\textrm{k}}^{\dagger}b_{-{\textrm{k}}}^{\dagger}\right)|0\rangle,$ (4)
with normalised Bogoliubov parameters $|\alpha_{\textrm{k}}|^{2}+|\beta_{\textrm{k}}|^{2}=1$. The density of produced fermions after the expansion is $n_{a}=n_{b}=\sum_{\textrm{k}\in{\rm BZ}}|\beta_{\textrm{k}}(\{\Gamma\}_{\eta})|^{2}/{{aN_{\rm s}\mathsf{a}_{\rm f}}}$, which is the result of a back-to-back production of particle-antiparticle pairs. We note that previous works [40, 41, 42, 43] on free QFTs have exploited the formal analogy of this transformation with two-mode squeezing to characterise mode entanglement in a cosmological scenario in momentum space, here connected to entanglement between right- and left-moving excitations. This, however, does not provide any spatial resolution that links to the causal propagation of entanglement, or to the existence of cosmological horizons due to an accelerated expansion. Moreover, to our knowledge, there has been no prior study of the effect of interactions and fermion condensates on the entanglement during gravitational expansion. We fill in both gaps in this work.
Correlation light cones.– Let us start moving away from the aforementioned mode entanglement by focusing on the block entanglement entropy (EE) $S_{A}=-\mathop{\rm Tr}(\rho_{A}\log\rho_{A})$, which quantifies the amount of quantum correlations across the $AB$ bipartition of the system $\rho_{A}=\mathop{\rm Tr}_{B}\ket{\psi}\!\!\bra{\psi}$ |, where $A$ would be a block of $\ell_{A}$ sites. We note that the EE has been studied in great detail for quantum quenches in many-body models, where a system evolves under a sudden change of the Hamiltonian, yielding a neat quasi-particle picture in which a linear entanglement growth is a manifestation of a light-cone-like propagation of excitations [44, 45, 46, 47, 48, 49, 50].
Given that the Hubble rate $\mathsf{H}$ allows us to smoothly interpolate between sudden and adiabatic limits, it is very natural to extend this EE quasi-particle picture to our cosmological scenario. Moreover, resorting to the concept of entanglement contour (EC) [28], which assigns to each site $i$ of the partition a specific contribution $S_{i}$ to the overall EE, such that $\sum_{i}S_{i}=S_{A}$, one can get further information on the local distribution of entanglement within the block, allowing us to bring this quasi-particle picture to a new domain. In particular, we show below that the EC allows identifying a cosmological particle horizon due to the accelerating spacetime. Although a general recipe for the calculation of EC in generic setups is lacking, the fGS ansatz provides a well-defined approximation to the EC via the aforementioned correlation matrix $\Gamma(\eta)$, as the spectrum of the block density matrix $\rho_{A}$ is determined by the restricted correlation matrix $(\Gamma_{A}(\eta))_{ij}=\Gamma_{ij}(\eta)|_{i,j\in A}$ [51, 52]. Diagonalizing this matrix by a unitary transformation $U^{\dagger}\Gamma_{A}U={\rm diag}(\{\nu_{\textrm{k}}\}_{\textrm{k}\in{\rm BZ}})$, the EC $S_{i}(\eta)=\sum_{\textrm{k}\in{\rm BZ}}p_{i}(\textrm{k},\eta)s_{\textrm{k}}(\eta)$ weighs each contribution $s_{\textrm{k}}(\eta)=-\nu_{\textrm{k}}(\eta)\log\nu_{\textrm{k}}(\eta)-(1-\nu_{\textrm{k}}(\eta))\log(1-\nu_{\textrm{k}}(\eta))$ at the $i$-th site by $p_{i}(k,\eta)=|U_{\textrm{k}i}(\eta)|^{2}$. Since this matrix is determined self-consistently (3), we can explore how entanglement evolves in the spacetime coordinates under interaction effects beyond perturbation theory, e.g. symmetry breaking. Let us start by studying how the particles created during the expansion propagate in the asymptotically-flat out region $\mathsf{a}_{\rm f}$, where the conformal time becomes linearly related to the cosmological time $\eta=\eta_{\rm f}+(t-t_{\rm f})/\mathsf{a}_{\rm f}$ and can attain arbitrarily-large values. In Fig. 1(a), we depict the spatiotemporal EC profile for a partition of $\ell_{A}=128$ sites in a $N_{\rm S}=512$ lattice in real time. In the free case $g_{0}^{2}=0$, one observes a clear light-cone-like propagation of entanglement. In the scaling limit, $\ell_{A}$, $t\to\infty$ with $t/\ell_{A}$ finite, this propagation of EC can be obtained by extending the EE quasi-particle picture to the EC, as discussed in the Supplemental Material [53]. The cones emanate from the block boundaries, separating causally connected and disconnected spacetime regions, and show a slope that is set by twice the group velocity $v_{\rm g}=\max_{\textrm{k}\in{\rm BZ}}\left|\partial_{\textrm{k}}\epsilon_{\textrm{k}}\right|$, where $\epsilon_{\textrm{k}}$ is the energy of a particle or antiparticle obtained by diagonalising $h_{\rm k}$ (1). Interestingly, this picture predicts that the entanglement equilibrates and there is an equipartition across the block $S_{i}(\eta)=S_{0}\,\,\forall i\in A$, such that the EE fulfills a volume law that agrees with a the eigenstate thermalization hypothesis [54, 55, 56] and a generalised Gibbs ensemble (GGE) that approximates $\rho_{A}$ [57].
When we switch on interactions to $g_{0}^{2}=1$, the scalar condensate attains a non-zero value in the initial state and displays a non-trivial time evolution $\Sigma(\eta)$. This results in a new dispersion relation from Eq. (3) that is corrected by a time-dependent self-energy and can, consequently, lead to a dynamical group velocity $v_{\rm g}(\eta)$. In this case, the previous conserved quantities associated to the GGE become time-dependent, and may compromise the validity of the quasi-particle prediction. However, for certain regimes in which the condensate relaxes to a certain value, one expects to be able to derive a hydrodynamic description for the EC by renormalising the spread velocity and the particle-antiparticle mode entanglement. We confirm this expectation in the right panel of Fig. 1(a), where we show that the EC light cones follow $x(\eta)\approx\pm 2{v}^{\rm int}_{g}\eta$, where ${v^{\rm int}_{g}}$ is a renormalised velocity. As shown in the figure, the cones get tilted with respect to $g_{0}^{2}=0$, showing that this propagation speed is enhanced by interactions. Although this is counterintuitive at first, as the continuum QFT predicts a dynamical mass generation that would slow the dynamics of excitations, the problem on the lattice is richer as this dynamically-generated mass can bring one closer or further apart from a critical line around which the continuum limit should be valid. In this case, increasing interactions brings us closer, reducing the gap and also the overall bare mass, such that the particle-anti-particle pairs propagate faster. Let us emphasise that this renormalised quasi-particle picture can actually be completely broken for stronger interactions in which the pseudo-scalar condensate is also non-vanishing $\Pi(\eta)$. We find in that case that the group velocity develops persistent oscillations that make a quasi-particle prediction very inaccurate EC (see the Supplemental Material [53]).
Correlation horizons.– Let us now turn to a key phenomenon that becomes apparent via the EC dynamics during the accelerated de Sitter expansion $\mathsf{a}(\eta)$, and after it as above. For this cosmological expansion, a particle horizon determines when the distance between the emission of a particle and its subsequent detection starts to expand faster than the speed of light. Accordingly, spacetime breaks in causally disconnected regions that can have no influence on one another and, thus, no entanglement. In conformal time, the appearance of this horizon in the EC is a bit subtle. In Fig. 1(b) we show the evolution of EC during the expansion, which shows a ballistic spread as before, but this time is upper bounded by $\eta=0$, as cosmological time squeezes in the distant future as a result of the exponential scale factor  [^1] To ensure significant propagation since the beginning of the expansion, we impose an initial state with matter content. This is achieved via a quantum quench in the mass before the start of the expansion. Other alternatives to design the initial state would lead to similar interplay with the horizon..
In Fig. 1(c), we show the EC profile in cosmological time $t$, which clearly shows how the propagation cones get curved as a consequence of the expansion of the universe, leading to a pair of asymptotes that that separate regions that cannot be causally connected. We see two separated spatial regions where entanglement attains non-vanishing values as time elapses but, due to the accelerated expansion, these are separated by a finite region with a width $\Delta x=\ell_{A}-\frac{4v_{g}}{{\sf H}\mathsf{a}_{0}}$ in which the EC is always vanishing. We emphasize that the spatial resolution brought by the EC as opposed to the EE is crucial to identify these correlation horizons. It is also interesting to note that the EC distribution within the causally-connected parts does not fulfill an equi-partition, so the evolution in this case is very different from an equilibration to a GGE. Had we let the system evolve in the Minkowski out region $\mathsf{a}_{\rm f}$, the particle-anti-particle pairs would have had sufficient time to propagate in this static background, such that the causally-disconnected regions would again be connected by subsequent correlation cones as before, leading to volume-law entanglement and GGE.
Symmetry breaking in the entanglement contour.– We have previously mentioned in passing that expansions in which the pseudo-scalar condensate is non-vanishing $\Pi(\eta)$ can have a bigger impact on the spread of entanglement. We now explore this further, considering that parity gets spontaneously broken, which will impact the properties of the EC under point-like transformations [28]. Specifically, the EC satisfies that $S_{i}=S_{j}$ if the system is invariant under a transformation that exchanges two sites, $\hat{U}_{ij}\hat{\psi}_{i}\hat{U}_{ij}^{-1}=\mathord{\rm e}^{\mathord{\rm i}\xi}\hat{\psi}_{j}$, where $\xi\in\mathbb{R}$ and $i\neq j$. To understand this property better, we temporarily focus on the discrete symmetries of the static model $\mathsf{a}(\eta)=1$, which connects to the tenfold classification of topological insulators [59, 60, 61]. For $\Pi=0$, the system is invariant under time-reversal $\hat{T}$ and particle-hole $\hat{C}$ transformations, as well as their combination $\hat{S}=\hat{T}\hat{C}$. At the single-particle level, these transformations are realised by $\mathsf{T}=T\mathcal{K}$ with $T=\gamma^{0}$, $\mathsf{C}=C\mathcal{K}$ with $C=\gamma^{0}\gamma^{1}$, and $\mathsf{S}=\gamma^{0}\gamma^{5}$, where $\mathcal{K}$ denotes complex conjugation such that $T^{\dagger}h_{-\textrm{k}}^{*}T=h_{\textrm{k}}$, $C^{\dagger}h_{-\textrm{k}}^{*}C=-h_{\textrm{k}}$ and $\mathsf{S}^{\dagger}h_{\textrm{k}}\mathsf{S}=-h_{\textrm{k}}$. These operators satisfy $\mathsf{T^{2}=C^{2}=S^{2}}=1$. In addition, parity symmetry $\hat{P}$ is realised as $P=\gamma^{0}$ such that $P^{\dagger}h_{-\textrm{k}}P=h_{\textrm{k}}$. Indeed, together with translational invariance, parity symmetry describes the invariance under a mirror symmetry about any point and, in particular, about the center of the $A$ block, which leads to a site exchange $\hat{U}_{ij}$ with $j=\ell_{A}+1-i$. According to our previous comment, if parity is conserved, one expects the EC to display the symmetry $S_{i}=S_{\ell_{A}+1-i}$, which agrees with the EC cones of Fig. 1.
This situation changes when $\Pi\neq 0$, as one finds that $\hat{T}$ is preserved, but $\hat{P}$, $\hat{C}$ and $\hat{S}$ get broken. Moreover, switching back to $\mathsf{a}(\eta)\neq 1$ will also break $\hat{T}$, endowing the condensate with its own dynamics $\Pi(\eta)$, which shall affect the EC spreading as we now discuss. In the left panel Fig. 2(a), we depict the spread of the EC in complete analogy to Fig. 1, but for an expansion with a non-zero $\Pi(\eta)$. We note that in all of the case above, the ECs displayed actually correspond to one of the spinor components, namely $S_{i}^{\rm u}(\eta)$, which does not compromise the generality as $S_{i}^{\rm d}(\eta)=S_{i}^{\rm u}(\eta)=:S_{i}(\eta)$. This no longer holds when $\Pi(\eta)\neq 0$, and one can observe in the right panel of Fig. 2(a) that the EC for the down spinor $S_{i}^{\rm d}(\eta)$ is indeed different, with a larger spread of entanglement from the left boundary with respect to the right one , opposite to what happens for $S_{i}^{\rm d}(\eta)$. These figures suggest that the EC profiles are still symmetric by simultaneously exchanging left-right boundaries and up-down spinors, yielding $S^{\textrm{u}}_{i}(\eta)=S^{\textrm{d}}_{\ell_{A}+1-i}(\eta)$ and viceversa $S^{\textrm{d}}_{i}(\eta)=S^{\textrm{u}}_{\ell_{A}+1-i}(\eta)$. In fact, one can check that this is indeed a remnant symmetry that combines both parity $\hat{P}$ and particle-hole $\hat{C}$ symmetries $(\gamma^{1})^{\dagger}h_{\textrm{k}}^{*}\gamma^{1}=-h_{\textrm{k}}$, which still holds even if the individual symmetries are broken. The ${C}{P}$ symmetry that survives a non-zero $\Pi(\eta)$ suggests that one can define a zigzag ordering of the spatial sites and spinor components and depict the site- and spinor-resolved EC as shown in Fig. 2(b). This symmetry-aware representation allows to recover a clear light-cone-like structure characteristic of causality and connecting with the aforementioned quasi-particle picture, even in presence of a pseudo-scalar condensate and for any Hubble rate $\mathsf{H}$.
Figure 2: Time evolution of EC for different phases for a partition $A$ of $N_{A}=32$ sites, from a system of $N_{S}=128$ sites. (a) Numerical time evolution of EC for each leg of the ladder, for an expansion with $\Pi\neq 0$, with $ma=-1$, $g^{2}=0$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, $\mathsf{H}=100$ (quench limit). (b) Numerical time evolution of EC, with all the sites shown in the horizontal axis, alternating up and down spinors, for an expansion with $\Pi\neq 0$, with $ma=-1$, $g^{2}=0$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, $\mathsf{H}=100$ (quench limit).
This situation contrasts the dependence on $\mathsf{H}$ of other more standard quantities in cQFTs, such as the spectrum $|\beta_{\textrm{k}}|^{2}$ of produced particles $n_{a}\sim\int{\rm d}\textrm{k}|\beta_{\textrm{k}}|^{2}/2\pi$. In general, the breakdown of parity translates into an asymmetric spectrum $|\beta_{\textrm{k}}|^{2}\neq|\beta_{-\textrm{k}}|^{2}$, which is natural in light of its effect $\hat{\psi}_{\textrm{k}}\rightarrow\hat{P}\hat{\psi}_{\textrm{k}}\hat{P}^{-1}=\gamma^{0}\hat{\psi}_{-\textrm{k}}$. However, as shown in Fig. 3, this is only manifested at intermediate values of the Hubble rate $\mathsf{H}$ in accordance with our previous results [38], but a symmetric spectrum around $\textrm{k}=0$ is recovered as one approaches the quench limit (see the darker blue curve). In the Supplemental Material [53], we indeed show how this peculiarity can be understood from the interplay of parity with the above discrete symmetries, applied to the instantaneous state after the sudden quench. We unveil an additional underlying symmetry that protects an equal particle production in modes ${\rm k}$ and $-{\rm k}$ regardless of the breaking of parity that depends on the adiabaticity parameter $\mathsf{H}$, as it is only present in the quench limit and not for slower expansions.
Figure 3: Spectra of particle production for various adiabaticity regimes. The parameters are set as $ma=-1$, $g_{0}^{2}=3$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, for a chain of $N_{\rm s}=128$ sites. We highlight four values of ${\sf H}a=4$, $0.3$, $0.2$ and $0.05$ for increasingly lighter blue, for which the relation between symmetry of the spectrum and $\mathsf{H}a$ is manifest.
Conclusions and outlook.– Our study highlights the potential for observing exotic analogues of gravitational particle production in effective cQFTs that go beyond the standard cosmological model. The possibility of tuning the dimensionality, designing the expansion scale factor, and tailoring the matter content of the cQFT brings a wide and largely-uncharted territory that can be addressed from a multidisciplinary viewpoint. In this work, we have focused on self-interacting Dirac QFTs in a reduced Friedman-Robertson-Walker spacetime, which have allowed us to explore how the typical phenomenon of particle-antiparticle production is affected by non-perturbative effects such as dynamical mass generation and the appearance of symmetry-broken fermion condensates. Exploiting a multidisciplinary perspective, we have analyzed this analogue cosmological setup through the lens of the entanglement contour. Making use of non-perturbative variational techniques based on fermionic Gaussian states, we have examined the spatio-temporal causal structure in the spread of entanglement carried by the produced particle-antiparticle pairs, and how the system can thermalise when looking at reduced parts. Our study reveals how the entanglement contour captures key phenomena, such as the cosmological particle horizon in a de Sitter spacetime, and how the effects of symmetry-breaking processes become very transparent when resolving the symmetries of the entanglement contour. Our work paves the way for future studies exploring further non-perturbative effects in interacting cQFTs, and encourages experimental efforts in Raman optical lattices to connect to these gravitational analogues.
Acknowledgements.– We acknowledge support from PID2021- 127726NB- I00 (MCIU/AEI/FEDER, UE), from the Grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033, from the the CAM/FEDER Project TEC-2024/COM 84 QUITEMAD-CM, and from the CSIC Research Platform on Quantum Technologies PTI- 001.